# Referee: Reference-aware Audiovisual Deepfake Detection

### v1

@@ 3.2 Implementation details and evaluation setup
Audiovisual encoder is initialized on Synchformer [25] pretrained on LRS3 [46]. Videos are sampled at 25 fps, while audio is sampled at 16 kHz and converted into 128-channel mel-spectrograms with 25 ms windows and a 10 ms hop. From each video, we extract a sequence of 8 overlapping segments with a duration of 0.64s (16 frames). The weighted sampling is adopted during training, while the network is optimized by Adam with 1e1e-55 learning rate. The learning rate is managed by a cosine annealing scheduler, which includes a linear warmup and decays to a minimum rate of 1e1e-66. During training, references are randomly sampled from real videos of the same identity, while evaluation uses predefined target–reference pairs. Importantly, the reference videos employed for training and testing are strictly disjoint to avoid data leakage. For evaluation, each video is divided into consecutive 2.88s windows with a 5% temporal overlap. The final video-level prediction is obtained by averaging the softmax probabilities across windows, followed by an argmax operation. Our code will be released upon acceptance.
Method Modality AUC AP
Xception [21] V 77.7 76.9
LipForensics [28] V 86.6 89.5
FTCN [29] V 68.1 66.8
RealForensics [31] V 93.6 95.7
AV-DFD [13] AV 82.1 79.6
AVAD [33] AV 86.9 87.6
AVFF [14] AV 95.5 93.1
Referee AV 97.5 98.1
Table 3: Cross-dataset performance comparison on KoDF.
Following previous works [12, 14], we assess the performance using accuracy (ACC), average precision (AP), and the area under the ROC curve (AUC). While ACC provides an intuitive measure of overall correctness but can be misleading under severe class imbalance, such as FakeAVCeleb, which contains 500 real and 21,066 fake videos. To address this issue, we additionally employ AP and AUC, both of which are threshold-independent. By using these three metrics, we can more reliably and comprehensively compare model performance, regardless of class imbalance within the dataset.
3.3 Results
Intra-dataset. Following prior works [16, 14], FakeAVCeleb is split 70% for training, and 30% for a test set. Table 1 shows that our method achieves superior improvements over visual-only and audiovisual baselines. Compared to RealForensics, the strongest visual model, our approach achieves gains of 9.3%p in ACC and 5.1%p in AUC, demonstrating the limitation of relying solely on visual cues. While audiovisual detectors such as AVFF already surpass uni-modal approaches, our Referee achieves SoTA with additional gains of 0.6%p in ACC, and 0.6%p in AUC. It demonstrates that, beyond simple multimodal fusion, explicitly modeling cross-modal biometrics consistency allows our approach to capture subtle discrepancies that current SoTA models miss.
Cross-dataset. We evaluate the generalization ability of our method in a cross-dataset setting, where the model is trained solely on FakeAVCeleb and tested on unseen datasets without any finetuning. As reported in Table 2, Referee outperforms all cross-dataset evaluated models in terms of both AUC and AP, demonstrating strong robustness to distribution shifts and unseen manipulation methods. Following the protocol from  [33, 14], Table 3 validates cross-lingual robustness, where our method achieves 97.5% AUC and 98.1% AP, surpassing previous SoTA results by more than 2%p in AUC and over 5%p in AP. It demonstrates that Referee shows consistent robustness to both zero-shot cross-lingual forgery and unseen manipulation methods. Overall, the results highlight the advantage of incorporating identity-aware representations, which enable our approach to mitigate the overfitting to dataset-specific artifacts observed in prior works.
@@ 3.3 Results
Referee 99.24 99.71 99.99
Table 4: Effectiveness of the identity matching design.
@@
Table 5: Performance varying the number of identity query tokens, and cross-attention layers.
Ablation study. Table 4 examines the effectiveness of the identity matching design. The removal of reference identity query leads to a noticeable drop in both ACC and AUC. This result highlights the necessity of our modeling reference-guided identity consistency. In contrast, when the identity matching module is retained but the identity loss is excluded, performance remains strong but slightly lower than the full model. This result indicates that our auxiliary matching task improves the robustness. These findings confirm that both the structural design of the ID matching module and its dedicated loss function are crucial for achieving optimal performance. Table 5 reports the performance varying the number of query tokens and cross-attention (CA) layers while keeping the IDB layers at two. Increasing the number of CA layers consistently improves performance across all metrics. While keeping the CA layers at two, expanding the number of query tokens from 4 to 6 yields a further performance gain, achieving the best overall results. While increasing the number of layers and queries further improves performance, we observe a clear trade-off with computational cost. This suggests that our chosen configuration strikes a favorable balance between accuracy and efficiency.
@@ 4 Conclusion
We introduce Referee, the reference-aware audiovisual framework that explicitly leverages a reference video for robust deepfake detection. Our identity bottleneck module implicitly learns cross-modal biometrics in ID queries for speaker verification. This identity-guided reasoning enables the model to move beyond superficial artifacts, offering stronger generalization to unseen manipulations and making it a resilient solution against the rapid progress of generative forgeries.

### v2

@@ 3.2 Implementation details
Audiovisual encoder is initialized on Synchformer [iashin2024synchformer] pretrained on LRS3 [afouras2018lrs3]. Videos are sampled at 25 fps, while audio is sampled at 16 kHz and converted into 128-channel mel-spectrograms with 25 ms windows and a 10 ms hop. From each video, we extract a sequence of 8 overlapping segments with a duration of 0.64s (16 frames). The model is optimized by Adam with an initial learning rate of 1e1e-55, following a cosine annealing scheduler with a linear warmup that decays to 1e1e-66. To account for class imbalance, weighted sampling is applied during training.
3.3 Evaluation protocols
For both training and evaluation, reference videos are sampled from real videos of the same identity as the target. Importantly, the real videos used for reference sampling are strictly disjoint between training and evaluation to ensure rigorous evaluation and prevent data leakage. During evaluation, each target video is partitioned into 2.88s windows with a 5% temporal overlap. A single 3s segment is randomly sampled from the reference video and consistently paired with all target windows. The final video-level prediction is determined by averaging the softmax probabilities across all windows.
Method Modality ACC AUC
Xception [rossler2019faceforensics++] V 67.9 70.5
LipForensics [haliassos2021lips] V 80.1 82.4
FTCN [zheng2021exploring] V 64.9 84.0
CViT [wodajo2021deepfake] V 69.7 71.8
RealForensics [haliassos2022leveraging] V 89.9 94.6
VFD [cheng2023voice] AV 81.5 86.1
AVoiD-DF [yang2023avoid] AV 83.7 89.2
AVFF [oorloff2024avff] AV 98.6 99.1
FRADE [nie2024frade] AV 98.6 99.8
CAD [du2025cad] AV 99.0 99.6
FoVB [nie2025towards] AV 98.5 99.7
Referee AV 99.2 99.7
Table 4: Effectiveness of the identity matching design.
@@ 3.3 Evaluation protocols
Following standard protocols [liu2024lips, oorloff2024avff], we report accuracy (ACC), average precision (AP), and the area under the ROC curve (AUC). Given the extreme class imbalance in datasets like FakeAVCeleb (500 real vs. 21,066 fake videos), we prioritize AUC and AP as they provide threshold-independent assessments of model robustness.
3.4 Results
Cross-dataset. We evaluate the generalization capability in a cross-dataset setting by testing on unseen datasets without any finetuning. As reported in Table 1, our Referee achieves the highest AP and competitive AUC, demonstrating that our model effectively generalized across unseen manipulation types and domain distribution shifts. Notably, while several reference-based approaches [poiforensics, idreveal] rely on the massive VoxCeleb2 [voxceleb2] dataset containing over 1 million utterances to train their identity encoders, Referee is trained using only 14k authentic clips. Despite this significant disparity in data scale, our framework effectively learns to capture robust biometric signatures, demonstrating that million-scale external datasets are not a prerequisite for learning universal identity priors. This resilience is further evidenced in the cross-lingual evaluation in Table 2, where our method surpasses previous SoTAs [nie2025towards, anshul2025intra] by achieving 99.41% AUC and 99.47% AP. Especially, we observe that the performance of existing reference-based models degrades significantly when encountering cross-lingual scenarios. By anchoring the detection process to cross-modal consistency rather than low-level pixel anomalies, our approach maintains high discriminative power even when encountering zero-shot linguistic contexts and unseen forgery techniques.
Intra-dataset. Table 3 shows that our method achieves superior improvements over visual-only and audiovisual baselines. Compared to RealForensics [haliassos2022leveraging], the strongest visual model, our approach achieves gains of 9.3%p in ACC and 5.1%p in AUC, demonstrating the limitation of relying solely on visual cues. Consistent with the superior performance of current audiovisual detectors [nie2025towards, du2025cad] over unimodal approaches, our Referee achieves the highest accuracy in intra-dataset evaluation. Explicitly modeling biometric consistency enables our approach to detect subtle discrepancies overlooked by existing fusion-based methods [nie2025towards, du2025cad].
Table 5: Performance varying the number of identity query tokens, and cross-attention layers in AV-Transformer.
@@ 3.4 Results
Ablation study. We note that all ablation studies were conducted on FakeAVCeleb [khalid2021fakeavceleb]. Table 4 examines the effectiveness of the core components in Referee. The removal of the reference identity query leads to a noticeable drop, as excluding identity loss results in a slight decline in robustness. These findings confirm that both our structural design and identity-aware supervision are essential for achieving optimal performance. Moreover, Table 5 reports the performance varying the number of query tokens and cross-attention (CA) layers in AV-Transformer while keeping the IDB layers at two. Increasing the number of CA layers improves performance across all metrics. While increasing the number of layers and queries further improves performance, we observe a clear trade-off with computational cost. This suggests that our chosen configuration strikes a favorable balance between accuracy and efficiency.
Analysis. Although FakeAVCeleb [khalid2021fakeavceleb] and FaceForensics++ [rossler2019faceforensics++] are constructed from YouTube videos and already incorporate real-world degradations like noise, compression artifacts, and varying resolutions, we further evaluate robustness under spatial and temporal perturbations to simulate practical scenarios. As shown in Table 6, performance remains stable even under 70% and 90% spatial cropping, indicating that the model does not rely on specific facial regions. Notably, Referee demonstrates remarkable stability to limited temporal information, achieving 99.71% AUC with just a 1-second reference clip. Overall, these results confirm that Referee is robust to spatial perturbations and variations in reference length, supporting its applicability in realistic deployment scenarios.
Table 6: Robustness analysis under spatial perturbation (random crop) and varying reference clip length on FakeAVCeleb.
Input noise ACC AUC AP
70% Random Crop 98.99 99.62 99.99
90% Random Crop 99.18 99.65 99.99
1 sec. Ref. clip 98.92 99.71 99.99
2 sec. Ref. clip 99.12 99.71 99.99
3 sec. Ref. clip 99.24 99.71 99.99
@@ 4 Conclusion
This study demonstrates that a reference-guided detection framework, Referee, provides a more generic solution for deepfake detection than conventional artifact-based approaches. Referee effectively identifies forgeries by detecting structural and biometric inconsistencies within the target stream. Our experiments show that this reference-to-target cross-modal modeling allows the model to remain invariant across different languages and manipulation types, achieving state-of-the-art results without relying on million-scale external datasets. Conclusively, by shifting the focus from ephemeral artifacts to explicit cross-modal feature correlation, our work establishes a robust and data-efficient defense against increasingly sophisticated digital manipulations.
5 Generative AI Use Disclosure
Generative AI tools were used only for improving language clarity and grammar. All scientific contributions and interpretations are entirely the work of the authors.